Femur and Humerus Fracture
K.B. is a 16-year-old, weighing 64 kg, who fell while skiing. She briefly lost consciousness but is now alert and oriented. She was transported down the hill by the ski patrol after being stabilized and then was flown to the hospital. She has a fractured right femur and humerus. She will be admitted to your unit after an open reduction and internal fixation (ORIF) of the femur fracture and casting of her leg and arm.
1. You are on the pediatric surgical floor and are taking report from the post anesthesia care unit (PACU) nurse. K.B. is awake and taking ice chips. For the orders listed in the chart below, state whether EACH order is appropriate or not and state why.
Chart View
Physician's Orders
1. Vital signs per routine
2. Neurologic checks every 4 hours
3. Turn, cough, and deep breathe and incentive spirometer every 2 hours while awake
4. Heat pack and elevate right lower extremity and right upper extremity
5. Neurovascular checks every 1 hour
6. NPO
7. IV fluids D51⁄2NS at 100 mL/hr
8. Morphine sulfate 5 mg IV every 4-6 hours prn
Case Study Progression:Case Study Progress
K.B. is settled into her room and begins to complain of pain (7 of 10) in her leg and arm. She weighs 65 kg. You note that the ordered dose of morphine sulfate was given 4 hours ago. Your drug reference states that the appropriate dose is 0.05 to 0.1 mg/kg every 4 to 6 hours.
2. Is this dose safe for your patient? Show your work.
3. The morphine for injection comes in a concentration of 2 mg/mL. How much will you draw up and have a second RN double-check?
Cast Study Progression:Case Study Progress
After K.B. has been on the unit for 6 hours, you identify the following changes in her assessment data:
K.B. is difficult to arouse, but when awake she is able to identify who and where she is; PERRLA 1+ with slower reaction time than earlier; color pale, pink; skin cool and clammy; heart rate 126 beats/min, respiratory rate 28 breaths/min, temperature (oral) 99 ° F (37.2 ° C); Spo2 90%. The findings of the neurovascular checks of the affected extremities are unchanged.
4. What will your immediate nursing interventions include? (List two possible interventions)
5. K.B.'s Glasgow Coma Scale score begins to decline from 15 to 11. What are possible reasons for changes in her neurologic status?
PLEASE ANSWER ALL QUESTIONS
Answer #1

1. 1) Vitals signs per routine is appropriate to detect any fluctuations in vital organ functioning.

2) Neurologic status checks every 4 hours , to see any deterioration in neurological functioning

3) Turn ,cough ,deep breaths and incentive spirometry every 2hour is appropriate and is taught pre surgery to expand lungs to prevent atelectasis and for doing it post surgery to avoid accumulation of secretion and pulmonary infections and hence prevent atelectasis.

4) Heat pack and elevate right lower extremity and right upper extremity is appropriate ,it is done to avoid inflammation and swelling

5) Neurovascular checks every 1 hour is appropriate to assess the neurovascular functioning

6) NPO to prevent aspiration patient is kept NPO before surgery

7) IV fluids D51⁄2NS at 100 mL/hr for hydration and fluid and electrolyte balance

8) Morphine sulfate 5 mg IV every 4-6 hours prn is appropriate and within normal range to relieve pain .

2. Weight of the patient is 65kg

Appropriate dose of morphine is 0.05mg to 0.1mg /kg every 4 to 6 hours .

No, the dose of morphine is not appropriate . For chronic pain we give 2.5 to 5mg /kg every 4hours in a child of 16 years .

3. The prescribed dose of morphine is 0.05- 0.1 mg /kg = 0.05×65 to 0.1×65 = 3.25 mg to 6.5mg

Available dose of morphine is 2mg/ml

Means 1ml has 2mg of morphine

We need 3.25mg to 6.5 mg

2mg of morphine is present in 1ml

1mg of morphine will be present in 1/2

Then 3.25mg of morphine will be present in 1/2 ×3.25 ml = 1.6ml

And 6.5mg of morphine will be present in 1/2× 6.5 ml = 3.2ml

1.6ml to 3.2ml of morphine will be drawn according to the upper and lower range and will be double checked by second RN .

4. Patient has sign and symptoms of hypotensive shock

2 Immediate interventions are :-

- Provide Trendelenburg position

- IV fluid administration (isotonic fluids like RL ,NS 0.9% etc )

5. The possible reasons for her changes in her neurologic status are :-

- shock ( hypotensive)

- Increased ICP